OpenStudy (anonymous):

The formula below gives A, the area of a parallelogram with base b and height h. A=bh If a parallelogram has an area of 60 square meters, which equation gives the height of the parallelogram as a function of its base? A) h=b/60 B)h=60b C)h=60-b D)h=60/b

jimthompson5910 (jim_thompson5910):

Start with A = bh and replace A with 60. We can do this because the area is 60 sq meters So we go from A = bh to 60 = bh

jimthompson5910 (jim_thompson5910):

Now how do we isolate h?

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

We have b times h on the right side, correct?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

We want to undo this to move that b over and isolate h. So what do we do to undo multiplication?

OpenStudy (anonymous):

divide?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so we divide both sides by b to isolate h to go from 60 = bh to 60/b = h

jimthompson5910 (jim_thompson5910):

The last step is to flip the equation to get h = 60/b

OpenStudy (anonymous):

so its b?

jimthompson5910 (jim_thompson5910):

what do you mean, it's h = 60/b

jimthompson5910 (jim_thompson5910):

or did you mean choice B?

jimthompson5910 (jim_thompson5910):

if you meant that, then no, it's not choice B

OpenStudy (anonymous):

i mean d

jimthompson5910 (jim_thompson5910):

yeah you got it

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw